Find three consecutive even integers such that five times the second equals twice the sum of the first and the third.
Answer by oscargut(686)

let n, n+2 and n+4 the numbers
then
5(n+2)=2(n+n+4)
5n+10=4n+8
then n=-2
so numbers are -2,0 and 2